OpenStudy (anonymous):

How does: ((sin(x))(1/x)-(ln(x))(cos(x))/(sin^2(x)) equal: ((sin(x))-(x)ln(x)(cos(x)))/((x)sin^2(x)) How are there more x's in the final answer?

zepdrix (zepdrix):

\[\Large \frac{\left(\frac{\sin x}{x}-\ln x \cos x\right)}{\sin^2x}\quad=\quad \frac{\left(\frac{\sin x}{x}-\frac{x \ln x \cos x}{x}\right)}{\sin^2x}\]

zepdrix (zepdrix):

Get a common denominator, see how that works? :)

zepdrix (zepdrix):

\[\Large =\frac{\left(\frac{\sin x-x \ln x \cos x}{x}\right)}{\sin^2x}\quad=\quad\frac{\sin x - x \ln x \cos x}{x \sin^2x}\]

OpenStudy (anonymous):

Thanks! :)

zepdrix (zepdrix):

c: